OpenStudy (anonymous):

The lengths of two sides of a triangle are shown below: Side 1: 3x2 - 2x - 1 Side 2: 9x + 2x2 - 3 The perimeter of the triangle is 5x3 + 4x2 - x - 3. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points)

OpenStudy (anonymous):

Part A: Rearrange the two equations so you can see the common terms together... side 1: 3x^2-2x-1 side 2: 2x^2+9x-3 So add the common terms together... 3x^2+2x^2=? -2x+9x=? -1+-3=?

OpenStudy (anonymous):

5x^2+7x=-4? @brookew41

OpenStudy (anonymous):

that equal sign doesn't belong in there... the sum of Side 1 and Side 2= 5x^2+7x-4. For part b all you have to do is the same thing, but subtract the sum of side 1 and 2 from the total. 5x^3-0x^3=? 4x^2-5x^2=? -x-7x=? -3-4=? As for part c, I have no idea what they are asking... Sorry!

OpenStudy (anonymous):

5x^3-x^2-8x-7 @brookew41

OpenStudy (anonymous):

??

OpenStudy (anonymous):

That's what I think.

OpenStudy (anonymous):

I know what part C means, it means that you need to explain how part A and B are related to the closure property. The closure property states that one polynomial plus another polynomial is also a polynomial. It is the same for addition, subtraction, multiplication, and division. So, on your module 7 quiz for FLVS's Algebra one, part two, you know what they are talking about.